Given that the kinetic energy of an object is 100 J. If the mass of object is 2kg then find its speed.
1
Expert's answer
2019-12-05T10:22:40-0500

From the formula for determining kinetic energy

"E_k=\frac{m \cdot v^2}{2}"

we write

"v^2=\frac{2 \cdot E_k}{m}"

then we get

"v=\sqrt{\frac{2 \cdot E_k}{m} }=\sqrt{\frac{2 \cdot 100}{2} }=\sqrt{100 }=10[m/s]"